# SAGE: Training Smart Any-Horizon Agents for Long Video Reasoning with Reinforcement Learning

### v1

@@ 1 Introduction
In the last year, there has been a natural shift from developing models for solely image reasoning [liu2023improvedllava, tong2024cambrian1, zhu2023minigpt, molmo2024, lin2023vila, jain2025visper_lm, jain2024vcoder, chen2023internvl, chen2024expanding, Qwen2-VL] to also tackling video reasoning [zhu2025internvl3, chen2024longvila, Qwen2.5-VL, wang2025internvl3_5, qwen3-vl, gemini25pushingfrontier] in the research community. Among the various model releases, the recent Gemini-2.5 [gemini25pushingfrontier] and Qwen3-VL [qwen3-vl] models pushed the frontier in video reasoning due to their ability to perform well on both short and long videos.
Although the aforementioned SOTA models differ in their training data, recipe, and architecture, among other things, they all function in a standard way when reasoning over videos: given a set of sampled frames, output the final answer with a single sequence prediction process, i.e., single turn reasoning. We refer to this line of work as falling under the Direct paradigm. Orthogonal to the works mentioned above, a few methods [liu2025videomind, wang2025videochat, zhi2025videoagent2enhancingllmbasedagent, chen2025lvagent, li2024wolf, goldfish] take an agentic route to predicting answers through multi-turn reasoning, falling under the Agent paradigm.
Humans excel at tasks that require multi-turn reasoning. For example, when viewing a 2-hour-long video, as humans, we take an iterative approach to finding the target information (Fig. 1). With the recent overwhelming success of RL post-training for training multi-turn agent systems for long-horizon tasks like software engineering [qwen3technicalreport, yang2024sweagent, faircodegenteam2025cwmopenweightsllmresearch], computer-use [yang2025gta1guitesttimescaling, seed2025seed1_5vl, qwen3-vl], and deep-research [dong2025tool, deepagent, tongyidr], it is natural to expect multi-turn agent systems to do well at long video reasoning. Despite the analogy above, most of the existing long video reasoning systems are still trained following the Direct paradigm, even with RL [chen2025longvila-r1, wang2025internvl3_5].
@@
(A2) Existing multi-turn agent systems usually use an LLM/VLM to orchestrate the calls to only a temporal grounder tool [liu2025videomind, videoexplorer, dang2025mupamultipathagenticreasoning] to iteratively locate an event over the entire video needed for finding an answer to a given question. However, we posit that attempting to ground an event in the whole video is not always the most effective approach due to the lack of robust temporal grounding models for long videos. For example, knowing the Formula 1 2024 season standings enables intelligent reasoning with a small temporal search space when watching the 2025 season livery reveal event video (Fig. 1a). Motivated by similar use cases, we introduce the SAGE (Smart Any-horizon aGEnt) system for long video reasoning. Particularly, we take a more innovative approach by equipping our system with tools such as web search and speech transcription, in addition to temporal grounding, to ensure that it is adept at not only utilizing visual signals from the video but also leveraging verbal and external knowledge. At the core of our system lies an orchestrator VLM, SAGE-MM, responsible for deciding between multi-turn and single-turn behavior for effective any-horizon reasoning. Moreover, guided by the fact that a user typically interacts with videos for entertainment [hafuta2025video, dannenbaum2025five], we focus our efforts on verifying the effectiveness of our approach on SAGE-Bench, curated with videos from popular YouTube channels to simulate use cases in the daily lives of users. Interestingly, we find existing agent systems to be over-engineered toward answering multiple-choice questions, often underperforming at the open-ended problems under SAGE-Bench (Tab. 4), demonstrating their ineffectiveness for real-world use-cases.
(A3) The variable duration of videos presents a unique challenge to training multi-turn agents. Specifically, during the RL post-training stage, the model should learn to function as an any-horizon agent, i.e., directly output the answer for simple problems while using multi-turn reasoning for harder problems [zhan2025katv1kwaiautothinktechnicalreport]. We believe that the optimization challenge posed by the dynamic nature of videos presents a challenge for training agent models using existing RL recipes, which have been shown to work well for training Direct models [chen2025longvila-r1, video-r1]. Moreover, extending the RLVR techniques [deepseekr1, deepseek-math] to video reasoning presents another challenge due to the task’s open-ended nature, which results in a lack of verifiable rewards. A few Direct approaches [wang2025video, chen2025longvila-r1] overcome the verifiable reward challenge by training only on MCQ problems and/or using some form of string-overlap metrics [video-r1, VideoRFT], rendering them ineffective at open-ended problems (Tab. 4).
@@
To that end, we propose a multi-reward RL recipe that utilizes strong reasoning LLMs [openai2024gpt4ocard] to validate the correctness of answers during the RL post-training stage. Moreover, moving away from using string-matching for evaluation, we adopt a universal LLM-as-a-judge evaluation approach to maintain uniformity across our training and evaluation setups. Our RL recipe improves the SFT model by 4.1% and surpasses the base by 5.7%, demonstrating its effectiveness. Moreover, for videos longer than 10 minutes, we observe performance improvements of up to 14.6% along with 4.8% for videos shorter than 10 minutes, proving SAGE’s effectiveness on any-horizon video reasoning.
@@ 2 Related Work
Table 1: Supported tools in SAGE. Our system has access to six tools, including web search (via the Serper-hosted Google Search API), for performing knowledge-driven reasoning. We implement the ground-event and analyze tools using existing MLLMs [qwen3-vl].
@@ 2.1 Long Video Reasoning Agents
VideoAgent [fan2025videoagent] creates a memory using the caption and keyframe features from the video subclips and incorporates tools to retrieve information from memory for reasoning. Similarly, VideoChat-A1 [wang2025videochat] employs keyframe retrieval to perform chain-of-shot reasoning. VideoMind [liu2025videomind] tunes LoRA adapters for the base Qwen2-VL [Qwen2-VL] model as a verifier to verify outputs from a separate temporal grounder module before final answer prediction. VideoExplorer [videoexplorer] optimizes the planner module with DPO [rafailov2023direct] for better trajectory reasoning. LVAgent [chen2025lvagent] leverages collaboration among multiple MLLMs with iterative reflection and key frame perception to reach the final answer.
@@ 2.2 Reinforcement Learning for Video Reasoning
Following the success of DeepSeek-R1 [deepseekr1] at using Reinforcement Learning with Verifiable Rewards (RLVR) to improve reasoning abilities in LLMs, various works have tried to leverage GRPO [deepseek-math, deepseekr1] to train Direct video reasoning models capable of thinking and then answering. Video-R1 [video-r1] follows the optimization approach of DeepSeek-R1 and introduces a contrastive temporal variant of GRPO, comparing answers between inputs with correct and incorrect frame ordering to enforce temporal dependence during reasoning. VideoRFT [VideoRFT] introduces a semantic-consistency reward between the reasoning trace and video frames. Video-Thinker [wang2025video] optimizes the model to output multiple temporal grounding instances within a single reasoning trace by carefully curating the cold-start SFT dataset. LongVILA-R1 [chen2025longvila-r1] enables the use of thousands of frames during the RL post-training stage with sequence parallelism. All the above methods utilize option-matching and ROUGE metrics to compute rewards, rendering their approach suboptimal for open-ended problems.
We train SAGE-MM to learn the ability to perform any-horizon reasoning using GRPO while leveraging an LLM-as-a-Judge to handle rewards for open-ended problems. A concurrent work, LongVT [yang2025longvt] also employs a similar training recipe with LLM-as-a-judge for computing the accuracy reward while only supporting a crop-video tool call.
@@ 3 Method
In the daily life of a human, entertainment is the primary purpose for interacting with videos [dannenbaum2025five, hafuta2025video], from watching sports videos on YouTube to scrolling through hundreds of short reels on Instagram. Therefore, it’s only natural to develop video reasoning models, keeping the user’s needs in mind. Among those needs, the open-ended interaction holds a vital place. For instance, as shown in Fig. 1, a user would usually ask: ”How does the Ferrari livery look this year?” as an open-ended question and expect the model to provide an answer in real-time. We introduce SAGE, a system designed to answer users’ questions while they enjoy entertainment videos. In the following subsections, we present technical details about SAGE (Sec. 3.1), followed by our synthetic data generation pipeline (Sec. 3.2). Lastly, we provide information on training the orchestrator (SAGE-MM) using RL for the system (Sec. 3.3).
@@ 3.1 System Design
The metadata string ($M$) comprises information about the video path and duration, which are necessary to predict the arguments for the tool call. We list the supported tools in SAGE in Tab. 1. Notably, unlike previous methods, which either perform temporal grounding over the complete video [liu2025videomind, videoexplorer], our SAGE autonomously predicts segment-level timestamps to ground events over a maximum duration of 10 minutes, as we qualitatively found that existing models struggle on longer entertainment videos.
@@ 3.2 Synthetic Data Generation
QnA Pairs. We leverage the long context modeling abilities of Gemini-2.5-Flash [gemini25pushingfrontier] to generate questions and answers for a given video in a single pass using a carefully designed prompt. We find that for videos longer than 5 minutes, having the model predict a percent_video_parsed field is critical to ensure that the generated questions temporally span the complete video, as shown at the bottom of Fig. 3. We generate 10-20 QnA pairs per video.
@@ 3.3 RL Post Training
We use GRPO [deepseek-math, deepseekr1] as the policy optimization algorithm during the RL post-training stage for trajectory-level optimization. Specifically, during the rollout generation, the $i^{th}$ action rollout trajectory for a given input set $S_{1}$ = $\{T,F,M,Q\}$ is represented by $\tau_{i}$. Therefore, we can formulate $\tau_{i}$ as a sequence of state-action pairs $\forall j\in[0,N]$:
@@
Accuracy Reward. We compute the outcome reward for a trajectory of length $N$ based on the final answer prediction using an LLM judge (GPT-4o [openai2024gpt4ocard]) to obtain a binary verdict indicating correctness at the last step.
@@
During training and inference, we set $N_{max}=11$ by default. However, during the RL stage, we find that setting $N_{max}=6$ for the first 100 steps is necessary for stable training, aligned with findings from a concurrent work for training long-horizon LLM agents [xi2025agentgymrltrainingllmagents]. Moreover, we penalize the model for predicting a wrong answer with tool calls to compensate for the positive step-level rewards while enforcing the any-horizon nature, i.e., making the model capable of predicting a direct answer. Conversely, we grant a slightly higher reward of +1.25 when the answer is correct and SAGE used visual tools (extract-video-parts or ground-event), reflecting the higher difficulty and importance of getting these tool calls right.
@@ 4 Experiments
For our experiments, we finetune for MLLMs, using both cold-start SFT (denoted by SFT) and RL post-training (denoted by RL) stages to obtain the SAGE-MM: Molmo-8B, Qwen2.5-VL-7B-Instruct [Qwen2.5-VL], Qwen3-VL-4B-Instruct [qwen3-vl], and Qwen3-VL-8B-Instruct [qwen3-vl]. During training, we freeze the visual encoder and projector modules. By default, we use the Qwen3-VL-8B-Instruct as the base SAGE-MM for all our ablations. We implement the transcribe-speech tool using the Whisper-large-v3 [radford2022whisper] model. We use the Qwen3-VL-30B-A3B-Instruct [qwen3-vl] model to perform temporal grounding and reasoning with the ground-event and analyze tools, respectively.
@@ 4.1 Implementation Details
Gemini-2.5-Flash [gemini25pushingfrontier] N/A Direct Direct 68.1 77.2 60.4 74.9 71.6 65.5
@@
GPT-4o [openai2024gpt4ocard] N/A Direct Direct 71.6 80.9 63.6 75.1 73.9 70.1
@@
Video-Thinker-7B [wang2025video] N/A Direct Direct 41.3 70.1 16.8 48.2 41.8 39.0
LongVILA-R1-7B [chen2025longvila-r1] N/A Direct Direct 52.6 68.8 38.7 57.6 64.9 49.6
VideoRFT-7B [VideoRFT] N/A Direct Direct 55.3 71.6 41.4 65.2 67.2 50.7
Video-R1-7B [video-r1] N/A Direct Direct 57.6 73.6 43.9 67.5 67.2 53.3
Qwen3-VL-30B-A3B-Instruct [qwen3-vl] N/A Direct Direct 67.6 81.3 55.8 72.8 71.6 65.4
VideoAgent [fan2025videoagent] GPT-4o N/A Agent 42.0 52.6 32.9 42.6 29.1 43.2
LVAgent [chen2025lvagent] InternVL-8/72B [chen2023internvl] + LLaVA-Video-72B [llavavideo] N/A Agent 49.7 70.5 32.1 54.1 48.5 48.4
LongVT [yang2025longvt] LongVT-7B-RFT Agent Agent 46.7 68.2 28.4 51.0 37.3 46.4
VideoMind [liu2025videomind] VideoMind-7B-Planner Agent Agent 50.0 69.7 33.2 50.8 41.8 50.7
VideoExplorer [videoexplorer] VideoExplorer-7B-Planner Agent Agent 50.1 69.6 35.1 52.0 40.2 51.3
VideoChat-R1.5 [yan2025videochatr15] VideoChat-R1.5-7B-M Agent Agent 54.8 73.8 38.6 55.1 48.5 55.1
Qwen2.5-VL-7B-Instruct [Qwen2.5-VL] N/A Direct Direct 58.6 74.2 45.4 65.8 68.7 55.2
@@
Qwen3-VL-4B-Instruct [qwen3-vl] N/A Direct Direct 62.7 75.8 51.6 69.3 66.4 60.2
@@
Qwen3-VL-8B-Instruct [qwen3-vl] N/A Direct Direct 64.9 77.7 54.0 72.8 68.7 61.9
@@
Molmo2-8B [molmo2] N/A Direct Direct 61.8 77.9 48.1 66.8 64.2 60.0
@@
Table 4: Comparison to Baselines. Using closed-source Gemini-2.5-Flash [gemini25pushingfrontier] and GPT-4o [openai2024gpt4ocard] as SAGE-MM improves upon the base models, showing the effectiveness of our system design. Our trained SAGE-MM also shows consistent improvements over all the baselines. SAGE-Flash refers to the setting where we use Gemini-2.5-Flash as the backend model for the ground-event and analyze tools. Existing Agent systems exhibit considerably worse performance on open-ended problems compared to our SAGE.
@@
Qwen2.5-VL-7B-Instruct [Qwen2.5-VL] N/A 32.7 37.8 25.8
VideoRFT-7B [VideoRFT] N/A 30.4 33.5 26.2
VideoMind-7B [liu2025videomind] N/A 30.7 34.0 26.2
Video-R1-7B [video-r1] N/A 31.5 36.0 25.8
VideoChat-R1.5-7B [yan2025videochatr15] N/A 33.8 35.3 31.8
@@
Table 5: Performance on MINERVA [minerva]. Our SAGE shows significant improvements on videos longer than 600 seconds.
@@
Evaluation. We evaluate all Direct baselines with 128 sampled frames as input, comparable to SAGE-MM’s input setting. Moreover, we also pass the video transcript as extra context to the Direct baselines for fair comparison. For Agent baselines, we follow their recommended setup. By default, we use LLM-as-judge (GPT-4o) for evaluating all models on both open-ended and MCQ problems. We set the temperature to 0.0 for all evaluations. However, because the action strings must follow a strict JSON schema, SAGE-MM occasionally produces malformed outputs. In such cases, we regenerate the response with a temperature of 0.7 for up to four attempts, which may lead to non-deterministic behavior during inference. We serve all supported models using vLLM [kwon2023efficient] during evaluation.

### v2

@@ 1 Introduction
In the last year, there has been a natural shift from developing models for solely image reasoning [29, 44, 63, 14, 28, 24, 23, 8, 7, 45] to also tackling video reasoning [64, 6, 2, 48, 42, 40] in the research community. Among the various model releases, the recent Gemini-2.5 [40] and Qwen3-VL [42] models pushed the frontier in video reasoning due to their ability to perform well on both short and long videos.
Although the aforementioned SOTA models differ in their training data, recipe, and architecture, among other things, they all function in a standard way when reasoning over videos: given a set of sampled frames, output the final answer with a single sequence prediction process, i.e., single turn reasoning. We refer to this line of work as falling under the Direct paradigm. Orthogonal to the works mentioned above, a few methods [30, 50, 61, 3, 26, 1] take an agentic route to predicting answers through multi-turn reasoning, falling under the Agent paradigm.
Humans excel at tasks that require multi-turn reasoning. For example, when viewing a 2-hour-long video, as humans, we take an iterative approach to finding the target information (Fig. 1). With the recent overwhelming success of RL post-training for training multi-turn agent systems for long-horizon tasks like software engineering [41, 54, 39], computer-use [55, 38, 42], and deep-research [16, 27, 43], it is natural to expect multi-turn agent systems to do well at long video reasoning. Despite the analogy above, most of the existing long video reasoning systems are still trained following the Direct paradigm, even with RL [5, 48].
@@
(A2) Existing multi-turn agent systems usually use an LLM/VLM to orchestrate the calls to only a temporal grounder tool [30, 57, 11] to iteratively locate an event over the entire video needed for finding an answer to a given question. However, we posit that attempting to ground an event in the whole video is not always the most effective approach due to the lack of robust temporal grounding models for long videos. For example, knowing the Formula 1 2024 season standings enables intelligent reasoning with a small temporal search space when watching the 2025 season livery reveal event video (Fig. 1a). Motivated by similar use cases, we introduce the SAGE (Smart Any-horizon aGEnt) system for long video reasoning. Particularly, we take a more innovative approach by equipping our system with tools such as web search and speech transcription, in addition to temporal grounding, to ensure that it is adept at not only utilizing visual signals from the video but also leveraging verbal and external knowledge. At the core of our system lies an orchestrator VLM, SAGE-MM, responsible for deciding between multi-turn and single-turn behavior for effective any-horizon reasoning. Moreover, guided by the fact that a user typically interacts with videos for entertainment [21, 12], we focus our efforts on verifying the effectiveness of our approach on SAGE-Bench, curated with videos from popular YouTube channels to simulate use cases in the daily lives of users. Interestingly, we find existing agent systems to be over-engineered toward answering multiple-choice questions, often underperforming at the open-ended problems under SAGE-Bench (Tab. 4), demonstrating their ineffectiveness for real-world use-cases.
(A3) The variable duration of videos presents a unique challenge to training multi-turn agents. Specifically, during the RL post-training stage, the model should learn to function as an any-horizon agent, i.e., directly output the answer for simple problems while using multi-turn reasoning for harder problems [58]. We believe that the optimization challenge posed by the dynamic nature of videos presents a challenge for training agent models using existing RL recipes, which have been shown to work well for training Direct models [5, 18]. Moreover, extending the RLVR techniques [13, 37] to video reasoning presents another challenge due to the task’s open-ended nature, which results in a lack of verifiable rewards. A few Direct approaches [47, 5] overcome the verifiable reward challenge by training only on MCQ problems and/or using some form of string-overlap metrics [18, 46], rendering them ineffective at open-ended problems (Tab. 4).
@@
To that end, we propose a multi-reward RL recipe that utilizes strong reasoning LLMs [33] to validate the correctness of answers during the RL post-training stage. Moreover, moving away from using string-matching for evaluation, we adopt a universal LLM-as-a-judge evaluation approach to maintain uniformity across our training and evaluation setups. Our RL recipe improves the SFT model by 4.1% and surpasses the base by 5.7%, demonstrating its effectiveness. Moreover, for videos longer than 10 minutes, we observe performance improvements of up to 14.6% along with 4.8% for videos shorter than 10 minutes, proving SAGE’s effectiveness on any-horizon video reasoning.
@@ 2 Related Work
Table 1: Supported tools in SAGE. Our system has access to six tools, including web search (via the Serper-hosted Google Search API), for performing knowledge-driven reasoning. We implement the ground-event and analyze tools using existing MLLMs [42].
@@ 2.1 Long Video Reasoning Agents
VideoAgent [17] creates a memory using the caption and keyframe features from the video subclips and incorporates tools to retrieve information from memory for reasoning. Similarly, VideoChat-A1 [50] employs keyframe retrieval to perform chain-of-shot reasoning. VideoMind [30] tunes LoRA adapters for the base Qwen2-VL [45] model as a verifier to verify outputs from a separate temporal grounder module before final answer prediction. VideoExplorer [57] optimizes the planner module with DPO [36] for better trajectory reasoning. LVAgent [3] leverages collaboration among multiple MLLMs with iterative reflection and key frame perception to reach the final answer.
@@ 2.2 Reinforcement Learning for Video Reasoning
Following the success of DeepSeek-R1 [13] at using Reinforcement Learning with Verifiable Rewards (RLVR) to improve reasoning abilities in LLMs, various works have tried to leverage GRPO [37, 13] to train Direct video reasoning models capable of thinking and then answering. Video-R1 [18] follows the optimization approach of DeepSeek-R1 and introduces a contrastive temporal variant of GRPO, comparing answers between inputs with correct and incorrect frame ordering to enforce temporal dependence during reasoning. VideoRFT [46] introduces a semantic-consistency reward between the reasoning trace and video frames. Video-Thinker [47] optimizes the model to output multiple temporal grounding instances within a single reasoning trace by carefully curating the cold-start SFT dataset. LongVILA-R1 [5] enables the use of thousands of frames during the RL post-training stage with sequence parallelism. All the above methods utilize option-matching and ROUGE metrics to compute rewards, rendering their approach suboptimal for open-ended problems.
We train SAGE-MM to learn the ability to perform any-horizon reasoning using GRPO while leveraging an LLM-as-a-Judge to handle rewards for open-ended problems. A concurrent work, LongVT [56] also employs a similar training recipe with LLM-as-a-judge for computing the accuracy reward while only supporting a crop-video tool call.
@@ 3 Method
In the daily life of a human, entertainment is the primary purpose for interacting with videos [12, 21], from watching sports videos on YouTube to scrolling through hundreds of short reels on Instagram. Therefore, it’s only natural to develop video reasoning models, keeping the user’s needs in mind. Among those needs, the open-ended interaction holds a vital place. For instance, as shown in Fig. 1, a user would usually ask: ”How does the Ferrari livery look this year?” as an open-ended question and expect the model to provide an answer in real-time. We introduce SAGE, a system designed to answer users’ questions while they enjoy entertainment videos. In the following subsections, we present technical details about SAGE (Sec. 3.1), followed by our synthetic data generation pipeline (Sec. 3.2). Lastly, we provide information on training the orchestrator (SAGE-MM) using RL for the system (Sec. 3.3).
@@ 3.1 System Design
The metadata string ($M$) comprises information about the video path and duration, which are necessary to predict the arguments for the tool call. We list the supported tools in SAGE in Tab. 1. Notably, unlike previous methods, which either perform temporal grounding over the complete video [30, 57], our SAGE autonomously predicts segment-level timestamps to ground events over a maximum duration of 10 minutes, as we qualitatively found that existing models struggle on longer entertainment videos.
@@ 3.2 Synthetic Data Generation
QnA Pairs. We leverage the long context modeling abilities of Gemini-2.5-Flash [40] to generate questions and answers for a given video in a single pass using a carefully designed prompt. We find that for videos longer than 5 minutes, having the model predict a percent_video_parsed field is critical to ensure that the generated questions temporally span the complete video, as shown at the bottom of Fig. 3. We generate 10-20 QnA pairs per video.
@@ 3.3 RL Post Training
We use GRPO [37, 13] as the policy optimization algorithm during the RL post-training stage for trajectory-level optimization. Specifically, during the rollout generation, the $i^{th}$ action rollout trajectory for a given input set $S_{1}$ = $\{T,F,M,Q\}$ is represented by $\tau_{i}$. Therefore, we can formulate $\tau_{i}$ as a sequence of state-action pairs $\forall j\in[0,N]$:
@@
Accuracy Reward. We compute the outcome reward for a trajectory of length $N$ based on the final answer prediction using an LLM judge (GPT-4o [33]) to obtain a binary verdict indicating correctness at the last step.
@@
During training and inference, we set $N_{max}=11$ by default. However, during the RL stage, we find that setting $N_{max}=6$ for the first 100 steps is necessary for stable training, aligned with findings from a concurrent work for training long-horizon LLM agents [52]. Moreover, we penalize the model for predicting a wrong answer with tool calls to compensate for the positive step-level rewards while enforcing the any-horizon nature, i.e., making the model capable of predicting a direct answer. Conversely, we grant a slightly higher reward of +1.25 when the answer is correct and SAGE used visual tools (extract-video-parts or ground-event), reflecting the higher difficulty and importance of getting these tool calls right.
@@ 4 Experiments
For our experiments, we finetune for MLLMs, using both cold-start SFT (denoted by SFT) and RL post-training (denoted by RL) stages to obtain the SAGE-MM: Molmo-8B, Qwen2.5-VL-7B-Instruct [2], Qwen3-VL-4B-Instruct [42], and Qwen3-VL-8B-Instruct [42]. During training, we freeze the visual encoder and projector modules. By default, we use the Qwen3-VL-8B-Instruct as the base SAGE-MM for all our ablations. We implement the transcribe-speech tool using the Whisper-large-v3 [35] model. We use the Qwen3-VL-30B-A3B-Instruct [42] model to perform temporal grounding and reasoning with the ground-event and analyze tools, respectively.
@@ 4.1 Implementation Details
Gemini-2.5-Flash [40] N/A Direct Direct 68.1 77.2 60.4 74.9 71.6 65.5
@@
GPT-4o [33] N/A Direct Direct 71.6 80.9 63.6 75.1 73.9 70.1
@@
Video-Thinker-7B [47] N/A Direct Direct 41.3 70.1 16.8 48.2 41.8 39.0
LongVILA-R1-7B [5] N/A Direct Direct 52.6 68.8 38.7 57.6 64.9 49.6
VideoRFT-7B [46] N/A Direct Direct 55.3 71.6 41.4 65.2 67.2 50.7
Video-R1-7B [18] N/A Direct Direct 57.6 73.6 43.9 67.5 67.2 53.3
Qwen3-VL-30B-A3B-Instruct [42] N/A Direct Direct 67.6 81.3 55.8 72.8 71.6 65.4
VideoAgent [17] GPT-4o N/A Agent 42.0 52.6 32.9 42.6 29.1 43.2
LVAgent [3] InternVL-8/72B [8] + LLaVA-Video-72B [59] N/A Agent 49.7 70.5 32.1 54.1 48.5 48.4
LongVT [56] LongVT-7B-RFT Agent Agent 46.7 68.2 28.4 51.0 37.3 46.4
VideoMind [30] VideoMind-7B-Planner Agent Agent 50.0 69.7 33.2 50.8 41.8 50.7
VideoExplorer [57] VideoExplorer-7B-Planner Agent Agent 50.1 69.6 35.1 52.0 40.2 51.3
VideoChat-R1.5 [53] VideoChat-R1.5-7B-M Agent Agent 54.8 73.8 38.6 55.1 48.5 55.1
Qwen2.5-VL-7B-Instruct [2] N/A Direct Direct 58.6 74.2 45.4 65.8 68.7 55.2
@@
Qwen3-VL-4B-Instruct [42] N/A Direct Direct 62.7 75.8 51.6 69.3 66.4 60.2
@@
Qwen3-VL-8B-Instruct [42] N/A Direct Direct 64.9 77.7 54.0 72.8 68.7 61.9
@@
Molmo2-8B [10] N/A Direct Direct 61.8 77.9 48.1 66.8 64.2 60.0
@@
Table 4: Comparison to Baselines. Using closed-source Gemini-2.5-Flash [40] and GPT-4o [33] as SAGE-MM improves upon the base models, showing the effectiveness of our system design. Our trained SAGE-MM also shows consistent improvements over all the baselines. SAGE-Flash refers to the setting where we use Gemini-2.5-Flash as the backend model for the ground-event and analyze tools. Existing Agent systems exhibit considerably worse performance on open-ended problems compared to our SAGE.
@@
Qwen2.5-VL-7B-Instruct [2] N/A 32.7 37.8 25.8
VideoRFT-7B [46] N/A 30.4 33.5 26.2
VideoMind-7B [30] N/A 30.7 34.0 26.2
Video-R1-7B [18] N/A 31.5 36.0 25.8
VideoChat-R1.5-7B [53] N/A 33.8 35.3 31.8
@@
Table 5: Performance on MINERVA [32]. Our SAGE shows significant improvements on videos longer than 600 seconds.
SAGE-$M$ Video-MMMU Video-MME
Qwen2.5-VL-7B-Instruct [2] N/A 57.5 63.6
Qwen3-VL-8B-Instruct [42] N/A 65.3 66.8
Video-R1-7B [18] N/A 61.5 61.2
SAGE-Flash (ours) Qwen3-VL-8B-Instruct [+SFT] 66.9 59.4
SAGE-Flash (ours) Qwen3-VL-8B-Instruct [+SFT] [+RL] 68.1 63.5
    - w/o ground-event 65.8 65.6
    - w/o ground-event & extract-video-parts 61.8 66.2
Table 6: Video-MMMU [22] & Video-MME [20] (w/o subs). Our SAGE-Flash outperforms baselines including Video-R1 [18] on Video-MMMU, demonstrating generalization to knowledge acquisition from videos.
@@
Evaluation. We evaluate all Direct baselines with 128 sampled frames as input, comparable to SAGE-MM’s input setting. Moreover, we also pass the video transcript as extra context to the Direct baselines for fair comparison. For Agent baselines, we follow their recommended setup. By default, we use LLM-as-judge (GPT-4o) for evaluating all models on both open-ended and MCQ problems. We set the temperature to 0.0 for all evaluations. However, because the action strings must follow a strict JSON schema, SAGE-MM occasionally produces malformed outputs. In such cases, we regenerate the response with a temperature of 0.7 for up to four attempts, which may lead to non-deterministic behavior during inference. We serve all supported models using vLLM [25] during evaluation.